OpenStudy (anonymous):

what would be 2x=-10

OpenStudy (kendricklamar2014):

Divide both sides by 2

OpenStudy (anonymous):

2x=-10 2/2x=-10/2 x=-5

OpenStudy (anonymous):

thanks

OpenStudy (kendricklamar2014):

Don't give out answers @AdmiralAngel25

OpenStudy (anonymous):

sorry...

OpenStudy (anonymous):

its ok at least she worked it out step by step

OpenStudy (anonymous):

he* lol

OpenStudy (kendricklamar2014):

Yeah, but if u look at the Code of Conduct, u have to lead them to the answer, not give the answer.

OpenStudy (leader):

2x=-10 -- -- 2 2 x=-5

OpenStudy (anonymous):

thanks